# In-plane Black-hole Spin Measurements Suggest Most Gravitational-wave Mergers Form in Triples

### v1

@@ 3 Results
to quantify the excess of near-perpendicular spin-orbit angles within $\pm\,\delta/2$ compared to near-aligned ($\cos\theta\in[1-\delta,1]$) and near-anti-aligned configurations ($\cos\theta\in[-1,-1+\delta]$), respectively. For the public hyper-parameter posterior samples of the B-Spline model (GWTC4pop; GWTC4pop-zenodo) and $\delta=0.1$ we find $Y_{\rm vs.~aligned}=1.5^{+0.7}_{-0.5}$ and $Y_{\rm vs.~anti-aligned}=2.5^{+1.8}_{-1.0}$, which suggest a statistically significant excess at $\cos\theta\approx 0$ and a skewness to positive values of $\cos\theta$.
Concerning our parametric models, we find that this spin distribution is best represented by Gaussian + Isotropic + Cut model which outperforms all other models, but Aligned + Isotropic + Cut, strongly (Jeffreys; Kass) with Bayes factors of $|\Delta\ln\mathcal{B}|>6.3$ (summarised in the last column of Table 1). This aligns with the astrophysical expectation of an isotropic component from dense environments above some mass threshold and a combination with mergers from stellar evolution processes below the threshold (e.g., Rodriguez2019; GerosaFishbach). Allowing for a free Gaussian is always statistically preferred over enforced spin-orbit alignment. Evidence of Gaussian + Isotropic + Cut against the aligned counterpart Aligned + Isotropic + Cut is positive but not decisive ($|\Delta\ln\mathcal{B}|=1.6$). However, in Appendix B we argue that the contribution from the aligned component in the Aligned + Isotropic + Cut needs to be small and gets more strongly disfavoured if very wide opening angles are excluded. Moreover, aligned models consistently disagree with the inferred excess fractions of the non-parametric model above, as summarised in Table 1. Thus, in what follows we focus on the Gaussian + Isotropic + Cut model. Other models are discussed at the end of this section and in Appendix B.
@@
The right panel shows excellent agreement with the non-parametric B-Spline model of GWTC4pop; GWTC4pop-zenodo exhibiting a similar peak near $\cos\theta\approx 0$, which lends further credibility to our analysis. The left panel shows that its Gaussian component, which is responsible for the overabundance at near-perpendicular orientations, agrees well with the expected tilt distribution (red) of mergers caused by the Lidov–Kozai effect in triples (Antonini2018). However, it is in stark contradiction with expectations from traditional isolated binary formation scenario (orange) which peak sharply at $\cos\theta=1$ (Olejak2024) even if large natal kicks are considered (dashed). We further discuss the astrophysical implications of our findings in Section 4.
@@
In Appendix A, we show that the PPD of the black hole spin tilt distribution of the low-mass Gaussian component is the result of posterior distributions with a small non-zero mean $\mu_{t}=0.20^{+0.21}_{-0.11}$ and standard deviation $\sigma_{t}=0.55^{+0.25}_{-0.16}$ which deviate significantly from their priors $\mu_{t}\sim\mathcal{U}_{[-1,1]}$ and $\sigma_{t}\sim\mathcal{N}_{[0.1,4]}(\mu=0,\sigma=1/2)$, respectively. We also find that the Gaussian component is the dominant component in the low-mass population ($\xi=0.86^{+0.10}_{-0.55}$) where the small isotropic contribution ($1-\xi$) causes the vertical shift of the PPD from the left to the right panel of Figure 2. For the mass threshold that separates the low-mass population from the high-mass isotropic population we infer $\tilde{m}=44.26^{+8.65}_{-4.57}\,\rm M_{\odot}$ which agrees with previous findings (2022ApJ...941L..39W; 2024PhRvL.133e1401L; 2025PhRvL.134a1401A; Antonini2025; Tong2025).
@@
In Figure 3, we present the PPDs for the spin magnitudes of the low-mass Gaussian component (parametrised by $\mu_{\chi}$ and $\sigma_{\chi}$), low-mass isotropic component ($\mu^{\rm Iso}_{\chi}$ and $\sigma^{\rm Iso}_{\chi}$), and high-mass isotropic component ($\mu^{\rm HighIso}_{\chi}$ and $\sigma^{\rm HighIso}_{\chi}$). The dominant low-mass Gaussian component (blue) tends to have small non-zero spins similar to findings of GWTC4pop. The subdominant low-mass isotropic component (orange) tends to exhibit somewhat larger spin magnitudes. However, the large uncertainty hinders reliable interpretations. Meanwhile, the high-mass isotropic component (purple) appears to peak at significantly larger values $\gtrsim 0.5$ which would be expected if those include highly spinning black holes formed through hierarchical mergers (GerosaFishbach) or through gas-accretion in active galactic nuclei. We highlight that the mismatches of each component with the non-parametric B-Spline model by the GWTC4pop; GWTC4pop-zenodo is expected as the latter fits for the entire merger population and does not differentiate between subpopulations.
@@
In Appendix B, we present the mixing fractions $\xi$ of the Gaussian component across different models (except Gaussian and Aligned where no mixing was assumed, cf. Table 1). The Aligned + Isotropic and Aligned + Isotropic + Cut models show that if alignment was enforced it would only contribute by $\xi=0.15^{+0.11}_{-0.06}$ and $0.09^{+0.66}_{-0.06}$, respectively, which further disfavours significant contribution from an aligned component. Comparing the Gaussian + Isotropic and Uniform Gaussian + Isotropic models to their LVK counterparts (which enforce the same spin magnitude distribution in the Gaussian and isotropic components) we observe a peculiar bi-modality around $\xi\approx 0.2$ and 0.8 in the former. However, the peak around 0.2 vanishes if only the mixing within the low-mass population in the Gaussian + Isotropic + Cut model is considered, which suggests that it is an artefact of the isotropic high-mass population.
@@
In Appendix B (Figure 8), we also show the PPDs of the spin-orbit tilts and spin magnitudes across all other models than Gaussian + Isotropic + Cut. In particular, it shows that models which enforce alignment of the Gaussian component (Aligned, Aligned + Isotropic, and Aligned + Isotropic + Cut) and Uniform Gaussian + Isotropic with a uniform prior on $\sigma_{t}$ fail to recover the peaked shape of the non-parametric B-Spline model (GWTC4pop; GWTC4pop-zenodo).
@@
Figure 3: PPD of the black hole spin magnitude distribution in the Gaussian Isotropic Cut model. Blue shows the magnitude distribution of the low-mass Gaussian component (parametrised by $\mu_{\chi}$ and $\sigma_{\chi}$), orange of the low-mass isotropic component ($\mu^{\rm Iso}_{\chi}$ and $\sigma^{\rm Iso}_{\chi}$), and purple of the high-mass isotropic component ($\mu^{\rm HighIso}_{\chi}$ and $\sigma^{\rm HighIso}_{\chi}$). The transition between low- and high-mass components is inferred at a primary mass cut-off of $\tilde{m}\approx 44.3^{+8.7}_{-4.6}\,\rm M_{\odot}$. The green colour shows the non-parametric B-Spline model of GWTC4pop; GWTC4pop-zenodo. Solid lines and shaded envelopes indicate medians and $90\,\%$ credible intervals, respectively.
4 Astrophysical implications
@@
Most massive progenitor stars of black holes or neutron stars are found in hierarchical triples or higher-order configurations (e.g., Moe2017; Offner2023), where a close inner binary is orbited by one or more outer distant companions. The gravitational three-body dynamics of hierarchical triples can naturally lead to binary mergers in which the component spins are nearly perpendicular to the inner orbital angular momentum (Antonini2018; Rodriguez2018). Liu2018 showed that this arises because the spin of a black hole in the inner binary evolves through a combination of de Sitter precession around the inner orbital angular momentum $\boldsymbol{L}\equiv\boldsymbol{L}_{\mathrm{in}}$, the precession of $\boldsymbol{L}_{\mathrm{in}}$ itself around the outer orbital axis $\boldsymbol{L}_{\mathrm{out}}$, and the gradual shrinking of the inner binary orbit due to gravitational-wave energy-loss.
@@
As the tertiary companion drives the inner binary to large eccentricities where efficient gravitational-wave emission at close pericentre passages shrinks the inner orbit, the coupling to the tertiary weakens, and $\Omega_{\rm SL}$ gradually overtakes $\Omega_{\rm L}$, causing $\boldsymbol{\hat{\Omega}}_{\rm p}$ to transition smoothly from being aligned with $\boldsymbol{L}_{\mathrm{out}}$ to being aligned with $\boldsymbol{L}_{\mathrm{in}}$. Because $\theta_{\rm p}$ is an adiabatic invariant, the spin maintains its inclination of $\simeq 90^{\circ}$ relative to the evolving precession axis. By the time the system decouples and the inner binary merges, the component spins therefore lie nearly within the orbital plane, leading to small values of $\chi_{\mathrm{eff}}$ but potentially large in-plane spin components. Meanwhile, the angles between the two spins in the orbital plane of the binary may be distributed across the whole range between 0 and $180^{\circ}$ due to generally different spin precession rates of the black holes if $m_{1}\neq m_{2}$ (Rodriguez2018, see Figure 10 therein).
@@
In contrast, binary black hole mergers formed through the evolution of isolated binary stars are generally expected to have spins closely aligned with the orbital angular momentum, resulting in low misalignment angles and a preference for positive effective spins (Kalogera2000; Farr2017; 2018PhRvD..98h4036G; Belczynski2020; Bavera2020; OlejakSpins2021; Broekgaarden2022). In standard binary-evolution models, stellar spins are assumed to be aligned with the orbital angular momentum prior to core collapse, and any natal kick imparted during the collapse tilts the orbital plane, thereby directly setting the resulting spin–orbit misalignment (e.g., Brandt1995; Kalogera2000; Fragos2010; Tauris2017), with the maximum tilt limited by the risk to disrupt the binary. As exemplified in Figure 2, even under the assumption of very high black hole natal kicks—which are not expected for massive black holes from an evolutionary perspective (Janka2024)—the resulting distribution remains strongly peaked at modest tilt angles, corresponding to $\cos\theta\gtrsim 0.75$.
@@
Producing a distribution with a significant fraction of systems at $\cos\theta\lesssim 0.75$ via isolated binary channels requires non-standard and poorly constrained assumptions, such as highly misaligned progenitor spins, “spin tossing” at black hole formation (Tauris2022), or finely tuned correlations between spin orientations and natal-kick directions (Baibhav2024). In particular, reproducing an observed peak near $\cos\theta\approx 0$ would require preferentially orienting black hole spins perpendicular to the natal kicks (Baibhav2024, see Figure 5 therein). Such an assumption is highly non-standard and is challenged by observational constraints from, e.g., pulsars’ velocity-spin alignment (Johnston2005; Noutsos2012; Noutsos2013; Mandel2023) as well as by simulation results of core-collapse supernovae (e.g., Burrows2024).
@@
Another possibility discussed in the literature is a spin flip induced by mass transfer (StegmannAntonini), which could produce in-plane tilts. This mechanism could reduce the first-born black hole’s spin projection onto the orbital angular momentum; however, its nature and efficiency remain highly uncertain as it requires very efficient angular momentum transport between the stellar core and envelope and inefficient tides which would otherwise realign the spins after mass transfer.
@@
Alternatively, VitaleMould2025 discuss the possibility that parametric modelling for an underlying preferentially aligned merger population may lead to inference of spurious peaks away from perfect alignment due to measurement uncertainties and the finiteness of the current sample size. We agree it is essential to verify the non-parametric inference of GWTC4pop and our parametric analysis with the growing gravitational-wave dataset in the future. However, we stress that VitaleMould2025 tested underlying preferentially aligned populations ($\mu_{t}=1$) which are extremely broad with $\sigma_{t}=1.15$, which seems inconsistent with current expectations from astrophysical models that tend to be narrower (see Section 2). Even in that case their spurious peaks away from alignment tend to be significantly higher than $\cos\theta\approx 0$.
5 Summary and discussion
@@
The latest gravitational-wave data GWTC-4.0 has enabled GWTC4pop to infer a global peak near $\cos\theta\approx 0$ in non-parametric models for the spin-orbit tilt angle distribution of the binary black hole merger population. In this work, we have recovered this distribution with parametric models that contain a dominant pronounced Gaussian peak at near-perpendicular directions. This defies traditional formation scenarios from isolated massive binary stars, which recover significant spin-orbit misalignment only under fine-tuned or highly uncertain assumptions of the binary evolution and stellar collapse (Section 4). Instead, the latest gravitational-wave data suggests an alternative explanation for the main origin of binary black hole mergers, which builds upon the fact that most massive black hole progenitor stars are found in hierarchical triple or higher-order configurations (Moe2017; Offner2023). In these systems, the relativistic gravitational dynamics of three-body systems naturally produces mergers with near-perpendicular spin angles (Antonini2018; Liu2018) and event rates consistent with gravitational-wave observations (MandelBroekgaarden2022, and references therein).
@@
We highlight that the spin distribution from triples also agrees qualitatively well with the inferred distribution of $\chi_{\rm eff}$, which is skewed and asymmetric about zero with more support for positive values (GWTC4pop; Banagiri2025b). Due to the natural preference to produce tilts at $\cos\theta\approx 0$ similar $\chi_{\rm eff}$ distributions have been obtained for triples, largely independent of the assumed spin magnitudes (e.g., Rodriguez2018, see Figure 9 therein). In addition, a triple formation scenario could contribute to the growing number of claims about mergers with residual orbital eccentricity (Romero-Shaw2022; Gupte2024; 2024ApJ...972...65I; 2025arXiv250415833D; Morras2025), which would be impossible to recover from isolated binary star evolution (e.g., Belczynski2002; Fumagalli2024). While some of the eccentric binary black hole candidates have masses that are confidently within the upper mass gap and may only plausibly explained by hierarchical mergers in dense environments (e.g., GW190521; Gayathri2020), candidates below the upper mass could be also explained by the Lidov–Kozai triples mechanism (2025PhRvD.112f3052R; McMillin2025). In particular, it has been argued that recent claims for residual eccentricity in one neutron star-black hole (NSBH) merger (Morras2025; planas2025; Kacanja2025; Jan2025) may only be obtained through triple star evolution (Stegmann2025) and imply a dominant contribution to the total NSBH merger rate (RomeroShawStegmannMorrasEccNSBH_inprep).
@@
If black hole mergers indeed tend to result from triple rather than binary star evolution, it would suggest that binary mass transfer is less efficient in forming very close binary black hole systems than previously assumed, e.g., in scenarios involving successful common envelope ejection (Belczynski2020). This may indicate that binary black hole progenitors instead tend to undergo stable mass transfer (Gallegos-Garcia2021), which typically produces systems with wider orbital separations than those expected from a common envelope scenario (Bavera2020; Olejak2021; vanSon2022). Moreover, recent studies have predicted that the minimum separation of binary black hole systems formed via binary evolution may be limited by a delayed unstable mass transfer and stellar mergers (Klencki2025). Also, observations of other types of binaries hosting compact objects further challenge standard binary evolution models. Notable examples are Gaia black holes (ElBadry2023a; Chakrabarti2023; ElBadry2023b), whose properties cannot be reconciled with conventional models (Nagarajan2025) and likely require non-standard assumptions about angular momentum loss during mass transfer (Olejak2025).

### v2

@@ III Results
to quantify the excess of near-perpendicular spin-orbit angles within $\pm\,\delta/2$ compared to near-aligned ($\cos\theta\in[1-\delta,1]$) and near-anti-aligned configurations ($\cos\theta\in[-1,-1+\delta]$), respectively. For the public hyper-parameter posterior samples of the B-Spline model (The LIGO, Virgo, and KAGRA Collaborations, 2025c, d) and $\delta=0.1$ we find $Y_{\rm vs.~aligned}=1.5^{+0.7}_{-0.5}$ and $Y_{\rm vs.~anti-aligned}=2.5^{+1.8}_{-1.0}$, which suggest a statistically significant excess at $\cos\theta\approx 0$ and a skewness to positive values of $\cos\theta$.
Concerning our parametric models, we find that this spin distribution is best represented by Gaussian + Isotropic + Cut model which outperforms all other models, but Aligned + Isotropic + Cut, strongly (Jeffreys, 1939; Kass and Raftery, 1995) with Bayes factors of $|\Delta\ln\mathcal{B}|>6.3$ (summarised in the last column of Table 1). Evidences for each model are calculated using the Dynesty nested sampler (Speagle, 2020), with error estimates on the order of $\mathcal{O}(0.1)$ in log-evidence, much smaller than the log Bayes factors we calculate in favour of the Gaussian + Isotropic + Cut model. This aligns with the astrophysical expectation of an isotropic component from dense environments above some mass threshold and a combination with mergers from stellar evolution processes below the threshold (e.g., Rodriguez et al., 2019; Gerosa and Fishbach, 2021). Allowing for a free Gaussian is always statistically preferred over enforced spin-orbit alignment. Evidence of Gaussian + Isotropic + Cut against the aligned counterpart Aligned + Isotropic + Cut is positive but not decisive ($|\Delta\ln\mathcal{B}|=1.6$). However, in Appendix C we argue that the contribution from the aligned component in the Aligned + Isotropic + Cut needs to be small and gets more strongly disfavoured if very wide opening angles are excluded. Moreover, aligned models consistently disagree with the inferred excess fractions of the non-parametric model above, as summarised in Table 1. Thus, in what follows we focus on the Gaussian + Isotropic + Cut model. Other models are discussed at the end of this section and in Appendix C.
@@
The right panel shows excellent agreement with the non-parametric B-Spline model of The LIGO, Virgo, and KAGRA Collaborations (2025c, d) and earlier parametric analyses of Vitale et al. (2022), which exhibit a similar peak near $\cos\theta\approx 0$ and lends further credibility to our analysis. The left panel shows that its Gaussian component, which is responsible for the overabundance at near-perpendicular orientations, agrees well with the expected tilt distribution (red) of mergers caused by the Lidov–Kozai effect in triples (Antonini et al., 2018). However, it is in stark contradiction with expectations from traditional isolated binary formation scenario (orange) which peak sharply at $\cos\theta=1$ (Olejak et al., 2024) even if large natal kicks are considered (dashed). We further discuss the astrophysical implications of our findings in Section IV.
@@
In Appendix B, we show that the PPD of the black hole spin tilt distribution of the low-mass Gaussian component is the result of posterior distributions with a small non-zero mean $\mu_{t}=0.20^{+0.21}_{-0.11}$ and standard deviation $\sigma_{t}=0.55^{+0.25}_{-0.16}$ which deviate significantly from their priors $\mu_{t}\sim\mathcal{U}_{[-1,1]}$ and $\sigma_{t}\sim\mathcal{N}_{[0.1,4]}(\mu=0,\sigma=1/2)$, respectively. We also find that the Gaussian component is the dominant component in the low-mass population ($\xi=0.86^{+0.11}_{-0.54}$) where the small isotropic contribution ($1-\xi$) causes the vertical shift of the PPD from the left to the right panel of Figure 2. A complete absence of the Gaussian component ($\xi=0$) is disfavoured but cannot be ruled out with certainty. For the mass threshold that separates the low-mass population from the high-mass isotropic population we infer $\tilde{m}=44.26^{+8.60}_{-4.56}\,\rm M_{\odot}$ which agrees with previous findings (Wang et al., 2022; Li et al., 2024; Antonini et al., 2025b, a; Tong et al., 2025; Plunkett et al., 2026).
@@
In Figure 3, we present the PPDs for the spin magnitudes of the low-mass Gaussian component (parametrised by $\mu_{\chi}$ and $\sigma_{\chi}$), low-mass isotropic component ($\mu^{\rm Iso}_{\chi}$ and $\sigma^{\rm Iso}_{\chi}$), and high-mass isotropic component ($\mu^{\rm HighIso}_{\chi}$ and $\sigma^{\rm HighIso}_{\chi}$). The dominant low-mass Gaussian component (blue) tends to have small non-zero spins similar to findings of The LIGO, Virgo, and KAGRA Collaborations (2025c). The subdominant low-mass isotropic component (orange) tends to exhibit somewhat larger spin magnitudes. However, the large uncertainty hinders reliable interpretations. Meanwhile, the high-mass isotropic component (purple) appears to peak at significantly larger values $\gtrsim 0.5$ which would be expected if those include highly spinning black holes formed through hierarchical mergers (Gerosa and Fishbach, 2021) or through gas-accretion in active galactic nuclei. We highlight that the mismatches of each component with the non-parametric B-Spline model by the The LIGO, Virgo, and KAGRA Collaborations (2025c, d) is expected as the latter fits for the entire merger population and does not differentiate between subpopulations.
@@
In Appendix C, we present the mixing fractions $\xi$ of the Gaussian component across different models (except Gaussian and Aligned where no mixing was assumed, cf. Table 1). The Aligned + Isotropic and Aligned + Isotropic + Cut models show that if alignment was enforced it would only contribute by $\xi=0.15^{+0.11}_{-0.06}$ and $0.09^{+0.66}_{-0.06}$, respectively, which further disfavours significant contribution from an aligned component. Comparing the Gaussian + Isotropic and Uniform Gaussian + Isotropic models to their LVK counterparts (which enforce the same spin magnitude distribution in the Gaussian and isotropic components) we observe a peculiar bi-modality around $\xi\approx 0.2$ and 0.8 in the former. However, the peak around 0.2 vanishes if only the mixing within the low-mass population in the Gaussian + Isotropic + Cut model is considered, which suggests that it is an artefact of the isotropic high-mass population.
@@
In Appendix C (Figure 10), we also show the PPDs of the spin-orbit tilts and spin magnitudes across all other models than Gaussian + Isotropic + Cut. In particular, it shows that models which enforce alignment of the Gaussian component (Aligned, Aligned + Isotropic, and Aligned + Isotropic + Cut) and Uniform Gaussian + Isotropic with a uniform prior on $\sigma_{t}$ fail to recover the peaked shape of the non-parametric B-Spline model (The LIGO, Virgo, and KAGRA Collaborations, 2025c, d).
@@
Figure 3: PPD of the black hole spin magnitude distribution in the Gaussian Isotropic Cut model. Blue shows the magnitude distribution of the low-mass Gaussian component (parametrised by $\mu_{\chi}$ and $\sigma_{\chi}$), orange of the low-mass isotropic component ($\mu^{\rm Iso}_{\chi}$ and $\sigma^{\rm Iso}_{\chi}$), and purple of the high-mass isotropic component ($\mu^{\rm HighIso}_{\chi}$ and $\sigma^{\rm HighIso}_{\chi}$). The transition between low- and high-mass components is inferred at a primary mass cut-off of $\tilde{m}\approx 44.3^{+8.7}_{-4.6}\,\rm M_{\odot}$. The green colour shows the non-parametric B-Spline model of The LIGO, Virgo, and KAGRA Collaborations (2025c, d). Solid lines and shaded envelopes indicate medians and $90\,\%$ credible intervals, respectively. The dotted lines show the prior bounds of the single-component models.
IV Astrophysical implications
@@
Most massive progenitor stars of black holes or neutron stars are found in hierarchical triples or higher-order configurations (e.g., Moe and Di Stefano, 2017; Offner et al., 2023), where a close inner binary is orbited by one or more outer distant companions. The gravitational three-body dynamics of hierarchical triples can naturally lead to binary mergers in which the component spins are preferentially nearly perpendicular to the inner orbital angular momentum (Antonini et al., 2018; Rodriguez and Antonini, 2018). Liu and Lai (2018) showed that this arises because the spin of a black hole in the inner binary evolves through a combination of de Sitter precession around the inner orbital angular momentum $\boldsymbol{L}\equiv\boldsymbol{L}_{\mathrm{in}}$, the precession of $\boldsymbol{L}_{\mathrm{in}}$ itself around the outer orbital axis $\boldsymbol{L}_{\mathrm{out}}$, and the gradual shrinking of the inner binary orbit due to gravitational-wave energy-loss.
@@
As the tertiary companion drives the inner binary to large eccentricities where efficient gravitational-wave emission at close pericentre passages shrinks the inner orbit, the coupling to the tertiary weakens, and $\Omega_{\rm SL}$ gradually overtakes $\Omega_{\rm L}$, causing $\boldsymbol{\hat{\Omega}}_{\rm p}$ to transition smoothly from being aligned with $\boldsymbol{L}_{\mathrm{out}}$ to being aligned with $\boldsymbol{L}_{\mathrm{in}}$. Because $\theta_{\rm p}$ is an adiabatic invariant, the spin maintains its inclination of $\simeq 90^{\circ}$ relative to the evolving precession axis. By the time the system decouples and the inner binary merges, the component spins therefore lie nearly within the orbital plane, leading to small values of $\chi_{\mathrm{eff}}$ but potentially large in-plane spin components. Meanwhile, the angles between the two spins in the orbital plane of the binary may be distributed across the whole range between 0 and $180^{\circ}$ due to generally different spin precession rates of the black holes if $m_{1}\neq m_{2}$ (Rodriguez and Antonini, 2018, see Figure 10 therein).
@@
In contrast, binary black hole mergers formed through the evolution of isolated binary stars are generally expected to have spins closely aligned with the orbital angular momentum, resulting in low misalignment angles and a preference for positive effective spins (Kalogera, 2000; Farr et al., 2017; Gerosa et al., 2018; Belczynski et al., 2020; Bavera et al., 2020; Olejak and Belczynski, 2021; Broekgaarden et al., 2022). In standard binary-evolution models, stellar spins are assumed to be aligned with the orbital angular momentum prior to core collapse, and any natal kick imparted during the collapse tilts the orbital plane, thereby directly setting the resulting spin–orbit misalignment (e.g., Brandt and Podsiadlowski, 1995; Kalogera, 2000; Fragos et al., 2010; Tauris et al., 2017), with the maximum tilt limited by the risk to disrupt the binary. As exemplified in Figure 2, even under the assumption of very high black hole natal kicks—which are not expected for massive black holes from an evolutionary perspective (Janka and Kresse, 2024)—the resulting distribution remains strongly peaked at modest tilt angles, corresponding to $\cos\theta\gtrsim 0.75$.
@@
Producing a distribution with a significant fraction of systems at $\cos\theta\lesssim 0.75$ via isolated binary channels requires non-standard and poorly constrained assumptions, such as highly misaligned progenitor spins, “spin tossing” at black hole formation (Tauris, 2022), or finely tuned correlations between spin orientations and natal-kick directions (Baibhav and Kalogera, 2024). In particular, reproducing an observed peak near $\cos\theta\approx 0$ would require preferentially orienting black hole spins perpendicular to the natal kicks (Baibhav and Kalogera, 2024, see Figure 5 therein). Such an assumption is highly non-standard and is challenged by observational constraints from, e.g., pulsars’ velocity-spin alignment (Johnston et al., 2005; Noutsos et al., 2012, 2013; Mandel and Igoshev, 2023) as well as by simulation results of core-collapse supernovae (e.g., Burrows et al., 2024).
@@
Another possibility discussed in the literature is a spin flip induced by mass transfer (Stegmann and Antonini, 2021), which could produce in-plane tilts. This mechanism could reduce the first-born black hole’s spin projection onto the orbital angular momentum; however, its nature and efficiency remain highly uncertain as it requires very efficient angular momentum transport between the stellar core and envelope and inefficient tides which would otherwise realign the spins after mass transfer.
@@
Alternatively, Vitale and Mould (2025) discuss the possibility that parametric modelling for an underlying preferentially aligned merger population may lead to inference of spurious peaks away from perfect alignment due to measurement uncertainties and the finiteness of the current sample size. This highlights that it is essential to verify the non-parametric inference of The LIGO, Virgo, and KAGRA Collaborations (2025c) and our parametric analysis with the growing gravitational-wave dataset in the future. However, we stress that Vitale and Mould (2025) tested underlying preferentially aligned populations ($\mu_{t}=1$) which are extremely broad with $\sigma_{t}=1.15$, which seems inconsistent with current expectations from astrophysical models that tend to be narrower (see Section II). Even in that case, their spurious peaks away from alignment are not as narrowly peaked as close to $\cos\theta\approx 0$ as the feature in the GWTC-4.0 data.
V Summary and discussion
@@
The latest gravitational-wave data GWTC-4.0 has enabled The LIGO, Virgo, and KAGRA Collaborations (2025c) to infer a global peak near $\cos\theta\approx 0$ in non-parametric models for the spin-orbit tilt angle distribution of the binary black hole merger population. In this work, we have recovered this distribution with parametric models that contain a dominant pronounced Gaussian peak at near-perpendicular directions. This defies traditional formation scenarios from isolated massive binary stars, which recover significant spin-orbit misalignment only under fine-tuned or highly uncertain assumptions of the binary evolution and stellar collapse (Section IV). Instead, the latest gravitational-wave data suggests an alternative explanation for the main origin of binary black hole mergers, which builds upon the fact that most massive black hole progenitor stars are found in hierarchical triple or higher-order configurations (Moe and Di Stefano, 2017; Offner et al., 2023). In these systems, the relativistic gravitational dynamics of three-body systems naturally produces mergers with near-perpendicular spin angles (Antonini et al., 2018; Liu and Lai, 2018), while they could also yield event rates (Mandel and Broekgaarden, 2022, and references therein) which are broadly consistent with gravitational-wave observations (The LIGO, Virgo, and KAGRA Collaborations, 2025c). Triples have also been shown to be in overall agreement with key features of the observationally inferred mass distributions such as a global peak at $\sim 10\,\rm M_{\odot}$ and declining rate towards higher masses (e.g., Antonini et al., 2017; Silsbee and Tremaine, 2017; Stegmann et al., 2022a, b). However, we note that there are no comprehensive studies that investigate the impact of assumptions in the stellar evolution (e.g., inner binary mass transfer or natal kicks) and account for delay time distribution and metallicity-dependent star-formation history, which are needed for robust comparison to gravitational-wave data.
@@
We highlight that the spin distribution from triples also agrees qualitatively well with the inferred distribution of $\chi_{\rm eff}$, which is skewed and asymmetric about zero with more support for positive values (The LIGO, Virgo, and KAGRA Collaborations, 2025c; Banagiri et al., 2025). Due to the natural preference to produce tilts at $\cos\theta\approx 0$ similar $\chi_{\rm eff}$ distributions have been obtained for triples, largely independent of the assumed spin magnitudes (e.g., Rodriguez and Antonini, 2018, see Figure 9 therein). In addition, a triple formation scenario could contribute to the growing number of claims about mergers with residual orbital eccentricity (Romero-Shaw et al., 2022; Gupte et al., 2024; Iglesias et al., 2024; de Lluc Planas et al., 2025b; Morras et al., 2025; Phukon et al., 2025), which would be impossible to recover from isolated binary star evolution (e.g., Belczynski et al., 2002; Fumagalli et al., 2024). While some of the eccentric binary black hole candidates have masses that are confidently within the upper mass gap and may only plausibly explained by hierarchical mergers in dense environments (e.g., Romero-Shaw et al., 2020; Gayathri et al., 2020), candidates below the upper mass could be also explained by the Lidov–Kozai triples mechanism (Romero-Shaw et al., 2025b; McMillin et al., 2025), where the subset of non-adiabatic triples is expected to yield a small fraction of eccentric outlier detections (e.g., Antonini et al., 2017; Rodriguez and Antonini, 2018; Liu et al., 2019). In particular, it has been argued that recent claims for residual eccentricity in one neutron star-black hole (NSBH) merger (Morras et al., 2025; de Lluc Planas et al., 2025a; Kacanja et al., 2025; Jan et al., 2025) may only be obtained through triple star evolution (Stegmann and Klencki, 2025) and imply a dominant contribution to the total NSBH merger rate (Romero-Shaw et al., 2025a).
@@
If black hole mergers indeed tend to result from triple rather than binary star evolution, it would suggest that binary mass transfer is less efficient in forming very close binary black hole systems than previously assumed, e.g., in scenarios involving successful common envelope ejection (Belczynski et al., 2020). This may indicate that binary black hole progenitors instead tend to undergo stable mass transfer (Gallegos-Garcia et al., 2021), which typically produces systems with wider orbital separations than those expected from a common envelope scenario (Bavera et al., 2020; Olejak et al., 2021; van Son et al., 2022). Moreover, recent studies have predicted that the minimum separation of binary black hole systems formed via binary evolution may be limited by a delayed unstable mass transfer and stellar mergers (Klencki et al., 2025). Also, observations of other types of binaries hosting compact objects further challenge standard binary evolution models. Notable examples are Gaia black holes (El-Badry et al., 2023b; Chakrabarti et al., 2023; El-Badry et al., 2023a), whose properties cannot be reconciled with conventional models (Nagarajan et al., 2025) and likely require non-standard assumptions about angular momentum loss during mass transfer (Olejak et al., 2025).